# Kolmogorov-Arnold Attention: Is Learnable Attention Better For Vision Transformers?

### v2

@@ 5.2 Performance Analysis
(i) Distribution of the Weights. [84] considered an invariant measure perspective to describe the training loss stabilization of the neural networks. We adopt this idea to study the distribution of the weights of the smallest and largest models of the ViTs, Tiny and Base, and their Fourier KArAt variants. Figure 3 shows the distribution of the weights of these models during different training phases—The evolution of the weights’ distributions for respective models remains invariant. Based on this observation, from [84], we can guarantee the convergence of loss values of all models; also, see Figure 7. However, as mentioned in [84], this perspective cannot comment on the generalization capacity and structural differences between different neural networks.
(ii) Loss Landscape. We visualize the loss landscape of KArAts to understand why they converge slowly in the later training phase, and why their generalizability does not scale with larger models. Following [45], we plot the loss surface along the directions in which the gradients converge; see §B.1.6 for implementation. These visualizations in Figure 4 show that Fourier KArAt in ViT architectures significantly impacts the smoothness of the loss surfaces. ViT-Tiny, with the fewest parameters, has the smoothest loss landscape and modest generalizability. In contrast, ViT-Tiny+Fourier KArAt’s loss landscape is spiky; it indicates the model is full of small-volume minima [35]. However, the model is modest in the number of parameters, so the gradient descent optimizer can still find an optimized model with better generalizability than the vanilla ViT-Tiny; hence, it gains a better test accuracy; see Table 2. ViT-Base, however, has more parameters than Tiny, and its loss surface is much spikier than ViT-Tiny. Finally, the loss surface of ViT-Base+Fourier KArAt is most spiky, making it a narrow margin model with sharp minima in which small perturbations in the parameter space lead to high misclassification due to their exponentially larger volume in high dimensional spaces; see Figure 4(d). Moreover, ViT-Base+$G_{3}B$ has 14 times more parameters than ViT-Tiny+$G_{1}B$. With learnable activations, gradient descent optimizers fail to find the best-optimized model, as small differences in the local minima translate to exponentially large disparities. The increasing number of local minima in the later phases of the training causes the slow convergence.
(iii) Attention Visualization. MHSA in vanilla ViTs captures region-to-region interaction in images. DINO [15] provides an innovative way to explain the dominant regions in an image that contribute towards the inference decision. It maps the self-attention values of the CLS token at the last encoder layer, which is directly responsible for capturing the information related to the class label. While the computer vision community predominantly uses this attention visualization, it is unsuitable for our case. Unlike softmax, Fourier KArAt does not inherently ensure learned attention values in $[0,1]$, as we dispense the $\ell_{1}$ projection. Therefore, we cannot interpret its performance similarly. Nonetheless, pre-softmax or pre-KArAt values in the attention matrix $\mathcal{A}^{i,j}$ are supposed to capture token-to-token interactions, and we adapt the attention maps to ignore the negative values for the sake of visualization. We use the ImageNet-1K trained models of ViT-Tiny (with traditional MHSA and Fourier KArAt) and plot the attention maps of the last layers in Figures 5 and 12.
@@
6 Discussion, Limitations, and Conclusion
@@
The learnable MHSA design raises many interesting questions regarding KArAt’s computing scalability and the resulting models’ generalizability. Below, we discuss their potential positive ($+$) and negative (--) implications and encourage researchers to elaborate on them.
@@
Designing KArAt self-attention is not parameter efficient and they are memory hungry (--). Having learnable activations in the MHSA causes a memory explosion on the present GPUs. Hence, calculations with the full-rank operator, ${\Phi\in\mathbb{R}^{N\times N}},$ are prohibitively expensive, regardless of the basis. Present GPUs and CUDA functions are not optimized for KAN implementation as noted by many before us; see [80, 66, 65, 78, 61]. Due to this, although parameter increase is not always significant (Table 2), optimized Fourier KArAt variants utilize 2.53×2.5-3$\times$2.5 - 3 × more GPU memory compared to traditional softmax attention. We propose a low-rank approximation to reduce memory requirements to a feasible level (Table 12), but researchers should explore alternative techniques. Interestingly, while there is a significant training time discrepancy between vanilla ViTs and Fourier-KArAts, their inference speeds are comparable on the present compute interface and hardware configuration; see Figure 8 in §B.2. Hence, we hope the future holds more efficient system-level optimization recipes for training KArAt. Additionally, we witnessed KArAts’ inconsistent training behavior with parameter changes in §B.3.4. Each ViT model with particular Fourier-KArAt variants has a typical $G$ value that brings out its best performance; there is no universal $G$ to follow.
@@
KArAts show a better attention score ($+$) and transfer learning capability ($+$) for all ViTs, but their generalizability does not scale with larger ViTs (--). Fourier KArAt concentrates high interaction scores on the object’s dominant features compared to having interactions throughout the object regions like vanilla ViTs [24, 73]. They exhibit decent transfer learning and generalizability. The distribution of weights guarantees that their training losses stabilize. Hence, by [84], we can postulate that first-order optimization algorithms (e.g., ADAM) can optimize KArAt’s loss landscape. But KArAt’s generalizability does not scale with larger models due to high parameterization, and [84] cannot explain this. We encourage researchers to investigate these models’ local or global optimality and generalizability through rigorous theoretical and empirical studies.
@@
Overparameterized KArAts’ weights lie in a much lower-dimensional subspace ($+$). Spiky loss landscape with local minima is probably the major impediment towards KArAt’s scalability; see Figure 4. Modern transformers are highly over-parameterized, and their weight matrices show low-rank structures [41]; spectral analysis (§B.5) of KArAt shows its attention matrices possess a better low-rankness than traditional attention (Figure 13). Considering this, properly guided parameter reduction in KArAt can result in an optimized model that generalizes well at scale.
@@
KArAts can provide better interpretability for smaller models ($+$). KANs are claimed to be more interpretable and accurate than MLPs due to their flexible choice of univariate functions [51]; it is also evident from KArAts’ attention maps. ViT-Tiny+KArAts outperforms vanilla ViT-Tiny on small-scale datasets, and performs at par on ImageNet-1K (Table 2) and transfer learning (Table 2). In the future, KArAt’s learnable activation makes a case for investigating it to add more explainability in vision tasks and can find more interpretability in smaller models trained in a limited data scenario.
@@
Conclusion.  Across our extensive benchmarking, KArAt outperforms traditional MHSA in smaller ViTs; results on larger models are mixed. Additionally, we note that the computing requirement impedes KArAt’s performance—training and fine-tuning these models are expensive. However, learnable activations can make a substantial difference in a model’s interpretability, and these limitations are not KArAt’s weaknesses. At its early stage, KArAt’s decent performance on diverse tasks shows its potential. In the future, one can extend this study beyond ViTs and check the resilience of KArAt in language processing and on large multimodal models [2, 79, 72, 13, 11, 3, 71, 4].
@@
Organization. We organized the Appendix as follows: §A contains the solution to problem (4). We also provide the pseudocode to implement Fourier-KArAt in ViT’s encoder block; see Algorithm 2. §B discusses additional numerical results, including implementation details, hyperparameter tuning, computation time analysis, diverse ablation studies, and different variants of Fourier KArAt. Finally, we concluded §B with the spectral analysis of the attention matrices for traditional and learnable attention.
Appendix A A Sparse Approximation Problem and its Solution
@@
$\begin{cases}&x^{\star}=\arg\min_{x\in\mathbb{R}^{m}}\frac{1}{2}\|x-y\|^{2}\\ &\text{subject to\;\;}x_{i}\geq 0\;\text{for}\;i\in[m],\text{and}\;\|x\|_{1}=1\text{.}\\ \end{cases}$ (4)
First, we rewrite the constrained problem (4) as an unconstrained problem using the Lagrange multipliers as:
@@
where $\lambda,\mu$ are Lagrange multipliers. Using the Karush-Kuhn-Tucker (KKT) stationarity condition on (5), we find $0\in\partial{\cal L}(x,\lambda,\mu)$, which implies, $x_{i}-y_{i}+\lambda=\mu_{i}.$ The complementary slackness gives, $\mu_{i}x_{i}=0$ for $i\in[m].$ Further, for $i\in[m]$, the primal and dual feasibility conditions are $x_{i}\geq 0,$ $\sum\limits_{i=1}^{m}x_{i}=1$, and $\mu_{i}\geq 0,$ respectively.
@@
Appendix B Addendum to Empirical Study
@@ B.1.1 Baselines and Datasets
B.1.2 Small-Scale Datasets used
@@
B.1.3 Implementing Different Basis and Base Activation Functions in KArAt
@@
For the experiments in the main paper, in the general form, (6), we use $b(x)=0$ for all $x\in\mathbb{R}$, grid size, $G=3$, and the row dimension of the ${\widehat{\Phi}^{i,j}}$ is set to $r=12$, or otherwise noted. We encourage the reader to explore other configurations that differ from those previously outlined.
(i) Different Basis Functions Used and Their Results. We incorporate a variety of widely used KAN activation functions into KArAt. By design, KArAt can utilize any basis function for activating the attention units. Therefore, in addition to the Fourier basis, we embed 5 different wavelet bases, and the Rational Function basis into KArAt. The function representations for these choices of bases are outlined in Table 6.
@@
(ii) Base Activation Functions Used and Their Results. We implement different base functions in ViT-Tiny+KArAt using the blockwise and universal configurations and show their performance on the image classification task on CIFAR-10; see the results in Table 8. For the majority of our experiments, including those in the main paper, we define and use ZeroModule and set the base activation function to be identically zero, or $b(x)=0$. Slightly lower computations within a single KAN layer help motivate us to use this base activation function. Overall, choosing between the activations SiLU and GELU for ViT-Tiny+KArAt with $G_{3}B,G_{3}U$ modes results in a slightly lower Top-1 accuracy compared to the results using the ZeroModule in the main paper (see Table 2).
Table 6: Different basis functions and their representations for $b(x)=0$. * For our experiments, the morlet central frequency hyperparameter is $\omega_{0}=5$, but other nonnegative values can be used. ** Meyer is defined to be $(m\circ\nu)(\tilde{x})=m(\nu(\tilde{x}))$, where $m(t)=\mathbb{I}\left(-\infty,\frac{1}{2}\right](t)+\mathbb{I}{\left(\frac{1}{2},1\right)}(t)\cos\left(\frac{\pi}{2}\nu(2t-1)\right)$ and $\nu(x)=x^{4}(35-84x+70x^{2}-20x^{3})\mathbb{I}{[0,1]}(x)$.
@@
DOG $-w\frac{d}{dx}\left(e^{-\frac{\tilde{x}^{2}}{2}}\right),\ \tilde{x}=\frac{x-\tau}{s}$
@@
Table 7: Blockwise and universal ViT-Tiny+KArAt and their Top-1 accuracies on CIFAR-10 using different base activation functions (nn.SiLU, nn.Identity, nn.GELU).
@@
B.1.4 Deploying Fourier KArAt to other ViTs
@@
(i) ConViT [20] uses two different attention mechanisms, the gated positional self-attention (GPSA) and MHSA. All of the variants of ConViT vary the number of attention heads per encoder while keeping the number of encoders the same. The purpose of the ConViT is to combine a CNN and a ViT in one encoder block. With the first ten layers using GPSA and the last two layers using MHSA, the model learns a balanced structure between CNNs and ViTs. Our variant of ConViT replaces the MHSAs with Fourier KArAts. We implement the universal and blockwise attention modes on this model. We give the ConViT’s hyperparameter configuration details in Table 9.
(ii) Swin Transformers [49] has a variable number of tokens. The number of tokens considered in the attention operation varies based on the window size, and thus, we do not implement a universal attention mode due to incompatibility. We replace regular and shifted window MHSAs, W-MHSA and SW-MHSA, with regular and shifted Fourier KArAts. We give the Swin-Transformer’s hyperparameter configuration details in Table 10.
B.1.5 Implementing Fourier KArAt for Object Detection and Semantic Segmentation

### v3

@@ 5.2 Performance Analysis
[…]
Attention Visualization. MHSA in vanilla ViTs captures region-to-region interaction in images. DINO caron2021dino provides an innovative way to explain the dominant regions in an image that contribute towards the inference decision. It maps the self-attention values of the CLS token at the last encoder layer, which is directly responsible for capturing the information related to the class label. While the computer vision community predominantly uses this attention visualization, it is unsuitable for our case. Unlike softmax, Fourier KArAt does not inherently ensure learned attention values in $[0,1]$, as we dispense the $\ell_{1}$ projection. Therefore, we cannot interpret its performance similarly. Nonetheless, pre-softmax or pre-KArAt values in the attention matrix $\mathcal{A}^{i,j}$ are supposed to capture token-to-token interactions, and we adapt the attention maps to ignore the negative values for the sake of visualization. We use the ImageNet-1K trained models of ViT-Tiny (with traditional MHSA and Fourier KArAt) and plot the attention maps of the last layers in Figures 4 and 11.
Figure 3: 3D-visualization of Loss landscape for ViT-Tiny and ViT-Base along the two largest principal component directions of the successive change of model parameters. KArAt’s loss landscapes are significantly less smooth than those of traditional attention; spiky loss landscapes are undesirable for optimization stability and the generalizability of the resulting model. See Figure 10 for the loss contours and the optimizer trajectory.
@@
5.3 KArAt’s Flexibility over Traditional Attention in ViT
Stemming from Kolmogorov-Arnold Representation Theorem (Theorem 1kolmogorov1956 , KANs liu2024kan with learnable activation functions can approximate complex mathematical functions and symbolic representations, and provide an interpretable alternative to MLPs liu2024kan ; liu2024kan2 ; yang2024kolmogorov . Precisely, if the target function admits a structure of the composition of $L$-KAN layers with smooth activation functions, then one can use an $L$-layer KAN with B-Spline basis to approximate it well [Theorem 2.1 in liu2024kan ]. This allows them to model complex relations that lie in the data distribution; KANs can discover the need for a new function whose numerical behavior suggests it may be a Bessel function (Figure 23(d) in liu2024kan ), and the authors also show that KANs can discover unknown equations (Figure 4(e)).
By using a similar argument, we hypothesize that KANs could be a better fit for approximating the underlying non-linear relationship between image patches and can be used for modeling token interactions with learnable components. Hence, KArAt can be viewed as an upgrade to vanilla self-attention, offering adaptability and more freedom for modeling token interactions in an interpretable way. KArAt provides the freedom to choose from a multitude of simple, easy-to-interpret, and tunable functions whose linear combination can produce a large class of unknown functions, rather than a fixed and known function, softmax. The traditional softmax attention adds a non-linearity for modeling such interactions in a one-size-fits-all mechanism without any justification. This interpretability is directly reflected in KArAt’s attention maps; see Figures 4 and 11.
Why KArAt’s attention maps are more explainable? DINO caron2021dino , being self-supervised, attends to the entire object and obtains richer features and deep visual representations ericsson2021well . To attend to the entire object like self-supervision, chefer2022optimizing in their supervised training, optimizes the attention map with additional loss objectives, which improves model interpretability and explanability. KArAt’s attention maps in Figures 4 and 11 capture entire objects without any extra loss, unlike some contours in the traditional softmax attention. It provides improved explanability and visual understanding, compared to spurious cues in supervised ViTs.
To show that KArAt’s attention maps are not just visually better, but the learnable attention leads to effective token interactions modeling, we follow the study by li2024surprising . To understand how the attention modulates the performance of ViTs, li2024surprising considered attention matrices from a masked auto-encoder he2022masked , trained with self-supervision, having superior performance, and used them in a vanilla ViT, ignoring the query and key, and training the rest of the network. This study shows that attention is the most important component in a transformer and can bring improved performance over the vanilla supervised ViT. In the same spirit, we use the precomputed attention matrices from the KArAt models to train the value branch of a vanilla ViT from scratch; key and query branches remain frozen. See the implementation details in §C.1 and the results in Table 4.
[…]
6 Discussion and Conclusion
@@
The learnable MHSA design raises many interesting questions regarding KArAt’s computing scalability and the resulting models’ generalizability. Below, we discuss their potential positive (++) and negative (-) implications and encourage researchers to elaborate on them.
@@
Designing KArAt self-attention is not parameter efficient and they are memory hungry (-). Having learnable activations in the MHSA causes a memory explosion on the present GPUs. Hence, calculations with the full-rank operator, $\Phi\in\mathbb{R}^{N\times N}$}, are prohibitively expensive, regardless of the basis. Present GPUs and CUDA functions are not optimized for KAN implementation as noted by many before us; see yang2024kolmogorov ; ruijters2012gpu ; ruijters2008efficient ; xu2024fourierkan ; KANsLimits . Due to this, although parameter increase is not always significant (Table 2), optimized Fourier KArAt variants utilize $2.5-3\times$ more GPU memory compared to traditional softmax attention. We propose a low-rank approximation to reduce memory requirements to a feasible level (Table 13), but researchers should explore alternative techniques. Interestingly, while there is a significant training time discrepancy between vanilla ViTs and Fourier-KArAts, their inference speeds are comparable on the present compute interface and hardware configuration; see Figure 7 in §C.3. Hence, we hope the future holds more efficient system-level optimization recipes for training KArAt. Additionally, we witnessed KArAts’ inconsistent training behavior with parameter changes in §C.4.4. Each ViT model with particular Fourier-KArAt variants has a typical $G$ value that brings out its best performance; there is no universal $G$ to follow.
@@
KArAts show a better attention score (++) and transfer learning capability (++) for all ViTs, but their generalizability does not scale with larger ViTs (-). Fourier KArAt concentrates high interaction scores on the entire object compared to spurious cues chefer2022optimizing like vanilla ViTs dosovitskiy2020vit ; deit . They exhibit decent transfer learning and generalizability. The distribution of weights guarantees the stability of the training. Hence, by zhang2022neural , we can postulate that first-order optimization algorithms (e.g., ADAM) can optimize KArAt’s loss landscape. But zhang2022neural cannot explain why KArAt’s generalizability does not scale with larger models due to high parameterization. We encourage researchers to investigate these models’ local or global optimality and generalizability through theoretical and empirical studies.
@@
Overparameterized KArAts’ weights lie in a much lower-dimensional subspace (++). Spiky loss landscape with local minima is probably the major impediment towards KArAt’s scalability; see Figure 3. Modern transformers are highly over-parameterized, and their weight matrices show low-rank structures kwon2024efficient ; spectral analysis (§C.7) of KArAt shows its attention matrices possess a better low-rankness than traditional attention (Figure 13). Considering this, properly guided parameter reduction in KArAt can result in an optimized model that generalizes well at scale.
@@
KArAts can provide better interpretability for smaller models (++). KANs are claimed to be more interpretable and accurate than MLPs due to their flexible choice of univariate functions liu2024kan . It is evident from KArAts’ attention maps, which are directly responsible for performance improvement; see §5.3. ViT-Tiny+KArAts outperforms vanilla ViT-Tiny on small-scale datasets, and performs at par on ImageNet-1K (Table 2) and transfer learning (Table 2). In the future, KArAt’s learnable activation makes a case for investigating it to add more explanability in vision tasks and can find more interpretability in smaller models trained in a limited data scenario.
@@
Conclusion.  Across our benchmarking, KArAt outperforms traditional MHSA in smaller ViTs; results on larger models are mixed. Additionally, we note that the computing requirement impedes KArAt’s performance—training and fine-tuning these models are expensive. However, learnable activations can make a substantial difference in a model’s interpretability, and these limitations are not KArAt’s weaknesses. At its early stage, KArAt’s decent performance on diverse tasks shows its potential. In the future, one can extend this study beyond ViTs and check the resilience of KArAt in language processing and on large multimodal models llama3 ; qwen2.5 ; thawakar2024mobillama ; campos2025gaea ; gpt3 ; llama32vision ; team2023gemini ; flamingo .
@@
Organization. We organized the Appendix as follows: §B contains the solution to problem equation 3. We also provide the pseudocode to implement Fourier-KArAt in ViT’s encoder block; see Algorithm 2. §C discusses additional numerical results, including implementation details, hyperparameter tuning, computation time analysis, diverse ablation studies, and different variants of Fourier KArAt. Finally, we concluded §C with the spectral analysis of the attention matrices for traditional and learnable attention.
Appendix A Multihead Self Attention—Continued
Appendix B A Sparse Approximation Problem and its Solution
@@
[…]
First, we rewrite the constrained problem equation 3 as an unconstrained problem using the Lagrange multipliers as:
@@
where $\lambda,\mu$ are Lagrange multipliers. Using the Karush-Kuhn-Tucker (KKT) stationarity condition on equation 4, we find $0\in\partial{\cal L}(x,\lambda,\mu)$, which implies, $x_{i}-y_{i}+\lambda=\mu_{i}.$ The complementary slackness gives, $\mu_{i}x_{i}=0$ for $i\in[m].$ Further, for $i\in[m]$, the primal and dual feasibility conditions are $x_{i}\geq 0,$ $\sum\limits_{i=1}^{m}x_{i}=1$, and $\mu_{i}\geq 0,$ respectively.
@@
Appendix C Addendum to Empirical Study
@@ C.1.1 Baselines and Datasets
C.1.2 Small-Scale Datasets used
@@
C.1.3 Implementing Different Basis and Base Activation Functions in KArAt
@@
For the experiments in the main paper, in the general form, equation 5, we use $b(x)=0$ for all $x\in\mathbb{R}$, grid size, $G=3$, and the row dimension of the $\widehat{\Phi}^{i,j}$} is set to $r=12$, or otherwise noted. We encourage the reader to explore other configurations that differ from those previously outlined.
Different Basis Functions Used and Their Results. We incorporate a variety of widely used KAN activation functions into KArAt. By design, KArAt can utilize any basis function for activating the attention units. Therefore, in addition to the Fourier basis, we embed 5 different wavelet bases, and the Rational Function basis into KArAt. The function representations for these choices of bases are outlined in Table 6.
@@
Base Activation Functions Used and Their Results. We implement different base functions in ViT-Tiny+KArAt using the blockwise and universal configurations and show their performance on the image classification task on CIFAR-10; see the results in Table 8. For the majority of our experiments, including those in the main paper, we define and use ZeroModule and set the base activation function to be identically zero, or $b(x)=0$. Slightly lower computations within a single KAN layer help motivate us to use this base activation function. Overall, choosing between the activations SiLU and GELU for ViT-Tiny+KArAt with $G_{3}B,G_{3}U$ modes results in a slightly lower Top-1 accuracy compared to the results using the ZeroModule in the main paper (see Table 2).
Table 6: Different basis functions and their representations for $b(x)=0$. * For our experiments, the morlet central frequency hyperparameter is $\omega_{0}=5$, but other nonnegative values can be used. ** Meyer is defined to be $)(m\circ\nu)(\tilde{x})=m(\nu(\tilde{x}$)), where $m(t)=\mathbb{I}\left(-\infty,\frac{1}{2}\right](t)+\mathbb{I}{\left(\frac{1}{2},1\right)}(t)\cos\left(\frac{\pi}{2}\nu(2t-1)\right)$ and $\nu(x)=x^{4}(35-84x+70x^{2}-20x^{3})\mathbb{I}{[0,1]}(x)$.
@@
DOG $w\frac{d}{dx}\left(e^{-\frac{\tilde{x}^{2}}{2}}\right),\ \tilde{x}=\frac{x-\tau}{s}$
@@
Table 7: Blockwise and universal ViT-Tiny+KArAt and their Top-1 accuracies on CIFAR-10 using different base activation functions (SiLU, Identity, GELU).
@@
C.1.4 Deploying Fourier KArAt to other ViTs
@@
ConViT [36] uses two different attention mechanisms, the gated positional self-attention (GPSA) and MHSA. All of the variants of ConViT vary the number of attention heads per encoder while keeping the number of encoders the same. The purpose of the ConViT is to combine a CNN and a ViT in one encoder block. With the first ten layers using GPSA and the last two layers using MHSA, the model learns a balanced structure between CNNs and ViTs. Our variant of ConViT replaces the MHSAs with Fourier KArAts. We implement the universal and blockwise attention modes on this model. We give the ConViT’s hyperparameter configuration details in Table 9.
Swin Transformers [37] has a variable number of tokens. The number of tokens considered in the attention operation varies based on the window size, and thus, we do not implement a universal attention mode due to incompatibility. We replace regular and shifted window MHSAs, W-MHSA and SW-MHSA, with regular and shifted Fourier KArAts. We give the Swin-Transformer’s hyperparameter configuration details in Table 10.
C.1.5 Implementing Fourier KArAt for Object Detection and Semantic Segmentation